OpenStudy (anonymous):

Check? Mrs. Stewart bought a rectangular plot of land. On a map, the coordinates of the vertices are (6, 0), (12, 3), (11, 5), and (5, 2). Each square unit on the map represents 4 acres. Mrs. Stewart purchased the land for $41,280. What was the price per acre? $ _________ 645?

OpenStudy (anonymous):

@Nnesha @Michele_Laino @bibby

OpenStudy (michele_laino):

we have to compute the base and the height of your rectangle

OpenStudy (michele_laino):

the base is given by the distance between the points (6,0) and (12,3): \[base = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} = \sqrt {{{\left( {12 - 6} \right)}^2} + {{\left( {3 - 0} \right)}^2}} = ...?\]

OpenStudy (anonymous):

√45?

OpenStudy (here_to_help15):

Still need help? 0.0

OpenStudy (anonymous):

yes please im not sure if @Michele_Laino is here anymore. @Here_to_Help15

OpenStudy (michele_laino):

ok! now the height of your triangle is the distance between teh points (5,2) and (6,0): \[height = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} = \sqrt {{{\left( {5 - 6} \right)}^2} + {{\left( {2 - 0} \right)}^2}} = ...?\]

OpenStudy (michele_laino):

hint: \[\begin{gathered} height = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} = \sqrt {{{\left( {5 - 6} \right)}^2} + {{\left( {2 - 0} \right)}^2}} = \hfill \\ = \sqrt {1 + 4} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

5 Sorry, i had to leave for a sec. c:

OpenStudy (here_to_help15):

lol ima ask again do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

we have: height=sqrt(5) so the area is: \[\Large area = \sqrt {45} \times \sqrt 5 = \sqrt {225} = ...?\]

OpenStudy (here_to_help15):

15 he just solved it for you : )

OpenStudy (michele_laino):

ok! :)

OpenStudy (michele_laino):

so we have numbers of acres = 15*4=...?

OpenStudy (michele_laino):

oops..number*

OpenStudy (anonymous):

60?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

So 60 is the answer?

OpenStudy (michele_laino):

no, since we have to compute the price per acre. Now that price per acre is: 41280/60=...?

OpenStudy (anonymous):

688?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thank you. c;

OpenStudy (michele_laino):

Thank you! :)